According to a Pew Research Center study, in May 2011, 37% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 338 community college students at random and finds that 135 of them have a smart phone. Then in testing the hypotheses:

H0: p = 0.37 versus

Ha: p > 0.37,

what is the test statistic?

z =  . (Please round your answer to two decimal places.)
Answer #1

solution

р = x / n = 135 / 338=0.399

P0 = 0.37

1 - P0 =1 - 0.37=0.63

Test statistic = z

= р - P0 / [\sqrtP0 * (1 - P0 ) / n]

=0.399 - 0.37 / [\sqrt(0.37*0.63) / 338] 0.02626

z= 1.10